Sarah Drogo, president of Storage, Inc. a company that makes a wide variety of storage boxes for home and office use, is thinking about adding a new line of small plastic storage boxes. This would require a new technology. The company currently uses predominantly cardboard of various weights that are used in its other products. Since this is a big move for the company, Sarah wants to make sure that all of the financial and nonfinancial implications are understood before she gets under way. She thinks she can sell 13 million units.

                Working with Jarrod White, her business analyst, Sarah puts together the following first pass information on the new boxes. They would retail for $1.99. The company sells to the distribution channel at 50% of retail, so the product would net the company approximately $1 per box. Jarrod estimates that the boxes would cost $0.15 in direct materials, $0.20 in direct labor, $0.05 in packaging materials, and $0.20 in variable machining costs.   The cost of the new equipment insurance, space (new space would be needed for the operation), support activities, and so on are estimated at $500,000 per year. In addition to these simple estimates, Jarrod determines that manufacturing the new box will require an additional person in manufacturing, quality control, and purchasing.   salaries would be $30,000, $35000, and $45,000, respectively.

                Adding the new line would also stress the packaging area. Jarrod believes there will be a onetime cost of $175,000 to retrofit the packaging and shipping area to accommodate the new smaller boxes, which will be packed in boxes of 12 for shipment to customers. The company currently bundle-packs its products with a simple strap device. The new containers will require their own cardboard box, which mean there will be many new activities in the packaging area. It is likely that to handle the increased workload, the company will need either a new person at a salary of $25,000 a year, or to pay overtime as $12.50 per hour for 10 hours a week to four existing workers.

                The new machine’s space requirements and noise level will be a challenge. Storage, Inc.’s existing machines are relatively quiet, but the high-pressure extrusion process needed for plastic is much nosier. It is also a bit more dangerous because of the high pressure required. Boxes can be damaged during the process, resulting in very sharp edges. Workers staffing the line will need to wear protective clothing and heavy gloves. Since, the plant is not air conditioned the combination of the heat from the machine and the heavy protective gear could make the job very uncomfortable. Sarah is considering air conditioning the part of the plant housing the new machine. This would drive the annual costs up from $500,000 to $650,000 a year, with a one-time cost of $75,000 for the sir conditioning units. Since it will be the only air-conditioned part of the plant, Sarah is concerned this will disgruntle the rest of the workforce.

                Full of questions, Sarah is looking for help to sort out the quantitative and qualitative issues involved in this potential expansion.

Required:

  1. Using as much of the information as possible, do a CVP analysis of the ongoing costs of running the new production line. How much does Sarah need to sell to breakeven? To make a $250,000 profit before tax? A $250,000 profit after tax with a tax rate of 25%? Do the analysis in two stages –Year 1 results with the nonrecurring expenditures and an ongoing production analysis for Year 2 that is believed to be the steady state for the line.
  2. Sarah could buy the new box from outside vendor in China for $0.80 per box. She would need new storage space to house the 250,000-unit minimum order required. This space would cost $30,000 a year. She should still need the new inspector to maintain quality control and the new shipping and packing capabilities as the order will be shipped in very large quantities bulked-packed. Review the numbers and do an incremental analysis of the make vs buy option Sarah faces.
  3. What should Sarah do? Use the numbers to justify your answer.
  4. What qualitative issues should concern Sarah? How much impact do you feel these should have on her decision? Why?
Answer #1
Variable costs per unit:
Manufacturing:
Direct materials $                   0.15
Direct labor $                   0.20
packaging material $                   0.05
Variable machining cost $                   0.20
Fixed costs per year:
Equipment insurance $      650,000.00
salaries $      105,000.00
packaging person salary $        25,000.00
(Since the number of weeks company operates is not given, 50 weeks working is assumed, thus the salary cost would be same under both the options overtime and new hiring)
Retrofitting cost $      175,000.00
Air conditioning cost $        75,000.00
TOTAL FIXED COSTS $ 1,030,000.00
1. BREAK EVEN POINT
FIXED COST / CONTRIBUTION PER UNIT =1030000/0.40    2,575,000 UNITS
SALES PRICE $          1.00
LESS:VARIABLE EXPENSES
Direct materials $                   0.15
Direct labor $                   0.20
packaging material $                   0.05
Variable machining cost $                   0.20 $          0.60
CONTRIBUTION PER UNIT $          0.40
2. PROFIT OF $250,000
LET NUMBER OF UNITS SOLD BE x
0.40 x -1030000 = 250000 128
0.40 x = 1280000 3200000
x= 3,200,000 units
3. PROFIT AFTER TAX OF $250,000
LET NUMBER OF UNITS SOLD BE x
(0.40 x -1,030,000 ) * 0.75 = 250000
0.30 x = 250000+ 772500 772500
0.30 x = 1,022,500 1022500
x= 3,408,333.33 units 3408333.333
OR 3,408,334 units
Variable costs per unit:
Manufacturing:
Direct materials $                0.15
Direct labor $                0.20
packaging material $                0.05
Variable machining cost $                0.20
Fixed costs per year:
Equipment insurance $   650,000.00
salaries $   105,000.00
packaging person salary $      25,000.00
(Since the number of weeks company operates is not given, 50 weeks working is assumed, thus the salary cost would be same under both the options overtime and new hiring)
TOTAL FIXED COSTS $   780,000.00
1. BREAK EVEN POINT
FIXED COST / CONTRIBUTION PER UNIT = 780000/0.40               1,950,000 UNITS
SALES PRICE $                     1.00
LESS:VARIABLE EXPENSES
Direct materials $                0.15
Direct labor $                0.20
packaging material $                0.05
Variable machining cost $                0.20 $                     0.60
CONTRIBUTION PER UNIT $                     0.40
2. PROFIT OF $250,000
LET NUMBER OF UNITS SOLD BE x
0.40 x -780000 = 250000
0.40 x = 1030000
x= 2,575,000 units
3. PROFIT AFTER TAX OF $250,000
LET NUMBER OF UNITS SOLD BE x
(0.40 x -780000 ) * 0.75 = 250000
0.30 x = 250000+ 585000 585000
0.30 x = 835,000 835000
x= 2,783,333.33 units 2783333.333
OR 2,783,334 units
Variable costs per unit: $                   0.80
Purchasing:
Fixed costs per year:
space $        30,000.00
salaries $        35,000.00
packaging person salary $        25,000.00
(Since the number of weeks company operates is not given, 50 weeks working is assumed, thus the salary cost would be same under both the options overtime and new hiring)
Retrofitting cost $      175,000.00
TOTAL FIXED COSTS $      265,000.00
Saving in fixed cost $      765,000.00
103000 - 265000
Less:
Increase in purchase cost $        50,000.00
250000 * (0.8-0.6)
increase in profit $      715,000.00

If sarah choose the purchase alternative than she has to terminate most of her employees